OpenStudy (pedro):

If p =2 , what is the value of 3p+2-p2 ? Show your work.

OpenStudy (anonymous):

3(2) +2 - 2^2 6 + 2 - 4 8 - 4 =4

jhonyy9 (jhonyy9):

6+2-4=6-2=4

OpenStudy (anonymous):

3(2) +2 - 2(2) 6 + 2 - 4 8 - 4 =4